A copper microstrip transmission line is to be designed for a 100 Ω characteristic impedance. The substrate is 0.51 mm thick, 𝜀𝑟 = 2.2 and tan δ = 0.001.


At 5 GHz, find:


3.1.1. The guide wavelength on the line (in cm) and;


3.1.2. The total attenuation in dB/cm.

1
Expert's answer
2022-03-02T10:06:43-0500

"\lambda=\frac {c}{\nu\sqrt{e_r}}=8.4~cm,"

"A=\gamma R\frac L{\sin\phi}=64~\frac{dB}{cm}."